OpenStudy (anonymous):

A population of wolves in a county is represented by the equation P(t)=80(0.98)^t, where (t) is the number of years since 1998. Predict the number of wolves in the population in the year 2008.

OpenStudy (australopithecus):

remember the rule log(x^(t)) = (t)log(x)

OpenStudy (australopithecus):

Take the log of both sides of the equation

OpenStudy (australopithecus):

to isolate the t

OpenStudy (australopithecus):

P(t)=80(0.98)^t = y/80 = (0.98)^t = log(y/80) = log(0.98^(t)) = log(y/80) = (t)log(0.98) = log(y/80)/log(0.98) = t Seriously over view the rules of logarithms you are never going to escape them if you plan on taking math in the future also they are neat

OpenStudy (australopithecus):

It is how we did large calculations back in the day

OpenStudy (australopithecus):

look at this site and try the problems http://www.sosmath.com/algebra/logs/log4/log44/log44.html